OpenStudy (anonymous):

What is the difference? 9/4-1/7

OpenStudy (anonymous):

you need a common denominator for both fractions before you can subtract. 

OpenStudy (phi):

can you find the common denominator ?

OpenStudy (anonymous):

I don't know what can be the denominator

OpenStudy (phi):

there are a few ways. the simplest is to multiply the two bottoms 4*7=28 that always gives a common denominator, but sometimes it gives you a bigger number than it has to be.

OpenStudy (anonymous):

9/4-1/7 = 63 - 4 / 28 = 59 / 28

OpenStudy (anonymous):

pick the smallest one from both columns that is the same